# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

### v1

@@ C.5 Pseudo Code of Swab
3:  Calculate VLM’s class-level zero-shot image classification rankings $\{r_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ and class-level gap vectors $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$.
@@
11:  Use $\bm{\gamma}^{*}$ and $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the class-level vectors $\{\bm{\hat{g}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset. Add $\{\bm{\hat{g}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ to corresponding text data’s feature of $\tilde{D}_{\mathcal{T}}$ to get modified text data.
12:  Use open-source datasets’ modified text data and images $D_{\mathcal{S}}^{I}$ to train the ranker model $f_{m}$.
13:  Use the ranker model $f_{m}$ to predict VLMs’ rankings $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ based on the target datasets’ modified text data.
14:  Use $\bm{\gamma}^{*}$ and $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the class-level zero-shot image classification rankings $\{\bm{\hat{r}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset. Calculate the average rankings $\{\hat{r}^{2}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ of $\{\bm{\hat{r}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ for each VLM.
15:  Ensemble $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ and $\{\hat{r}^{2}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ to get final predicted rankings $\{\hat{r}^{ens}_{m,\mathcal{T}}\}_{m=1}^{m=M}$.
16:  Select the best VLM based on $\{\hat{r}^{ens}_{m,\mathcal{T}}\}_{m=1}^{m=M}$.
Appendix D More Experiment Results
@@
D.1 Bridging the Modality Gap Leads to Better Image Proxies
@@
We use ImageNet as our dataset. First, we employ the method introduced in subsection 3.2 to predict the gap vectors for each class of the target dataset based on gap vectors calculated on open-source datasets. Then, we add the corresponding class’s gap vectors to the generated text data used in Section 2.3 and Figure 2 to bridge the modality gap. Finally, we calculate the accuracy of different models on these modified text data. We also calculate the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better) to measure the consistency between the predicted Top-1 accuracy and the true image classification accuracy.
We compare the consistency metrics of text data and modified text data. It can be observed that the consistency metrics of modified text data are better, which proves our method can reduce the gap between the generated text data and the images.
D.2 Per-Dataset Experiment Result
@@
Table 9: LOVM Benchmark (top-1 accuracy).We compare the per-dataset experiment result between ModelGPT and Swab.
Stanford Cars CIFAR100 CLEVR-DIST. CLEVR-COUNT Country211 Retinopathy DMLab DTD EuroSAT FER2013 Flowers102 GTSRB ImageNet KITTI MNIST PCam Oxford Pets Rendered SST2 RESISC45 STL10 SUN397 SVHN VOC2007 Mean $R_{5}$ ModelGPT 0.80 0.80 0.00 0.60 0.60 0.00 0.00 1.00 0.60 0.20 0.60 0.60 0.80 0.00 0.20 0.00 0.80 0.00 1.00 0.20 0.60 0.40 0.60 0.452 Swab 0.40 0.60 0.00 0.60 0.40 0.20 0.00 0.80 0.60 0.20 0.60 0.40 0.80 0.00 0.60 0.20 0.80 0.60 0.80 0.20 0.80 0.60 0.80 0.478 $\tau$ ModelGPT 0.33 0.67 0.00 -0.33 -0.33 0.00 0.00 -0.20 1.00 0.00 1.00 1.00 1.00 0.00 0.00 0.00 0.33 0.00 -0.20 0.00 -1.00 0.00 1.00 0.186 Swab 0.00 1.00 0.00 0.33 0.00 0.00 0.00 0.00 1.00 0.00 1.00 0.00 0.33 0.00 0.33 0.00 0.33 1.00 0.00 0.00 0.33 0.33 0.33 0.275

### v3

@@ D.4 Pseudo Code of Swab
3:  Calculate VLM’s class-level zero-shot image classification rankings $\{r_{m,i}^{\mathcal{S}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ and class-level gap vectors $\{\bm{g}_{m,i}^{\mathcal{S}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$.
@@
11:  Use $\bm{\gamma}^{*}$ and $\{\bm{g}_{m,i}^{\mathcal{S}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the class-level vectors $\{\bm{\hat{g}}_{m,i}^{\mathcal{T}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset. Add $\{\bm{\hat{g}}_{m,i}^{\mathcal{T}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ to corresponding text data’s feature of $\tilde{D}_{\mathcal{T}}$ to get modified text data.
12:  Use open-source datasets’ modified text data and VLMs’ ground truth zero-shot image classification performance to train the ranker model $f_{m}$.
13:  Use the ranker model $f_{m}$ to predict VLMs’ rankings $\{\hat{r}^{\mathcal{T},(1)}_{m}\}_{m=1}^{m=M}$ on the target dataset based on the target dataset’s modified text data.
14:  Use $\bm{\gamma}^{*}$ and $\{\bm{r}_{m,i}^{\mathcal{S}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the VLMs’ class-level zero-shot image classification rankings $\{\bm{\hat{r}}_{m,i}^{\mathcal{T}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset.
15:  Calculate the average prediction rankings across all classes in the target dataset $\{\hat{r}^{\mathcal{T},(2)}_{m}\}_{m=1}^{m=M}$ of $\{\bm{\hat{r}}_{m,i}^{\mathcal{T}}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ for each VLM as the VLM’s overall prediction ranking on the target dataset.
16:  Ensemble $\{\hat{r}^{\mathcal{T},(1)}_{m}\}_{m=1}^{m=M}$ and $\{\hat{r}^{\mathcal{T},(2)}_{m}\}_{m=1}^{m=M}$ to get final predicted rankings $\{\hat{r}^{\mathcal{T},ens}_{m}\}_{m=1}^{m=M}$.
Appendix E More Experiment Results
@@
E.1 Bridging the Modality Gap Leads to Better Image Proxies
@@
We use ImageNet as our dataset. First, we employ the method introduced in Section 3.2 to predict the gap vectors for each class of the target dataset based on gap vectors calculated on open-source datasets. Then, we add the corresponding class’s predicted gap vectors to the generated text data of ImageNet to bridge the modality gap. Finally, we calculate the zero-shot classification accuracy of different models on these modified text data. To measure the consistency between the predicted Top-1 accuracy and the true image classification accuracy, we calculate the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better). We compare the consistency metrics of text data and modified text data. It can be observed that the consistency metrics of modified text data are better, which proves our method can reduce the gap between the generated text data and the image data.
Figure 4: Comparison of the consistency metrics between the accuracy calculated using text data before and after bridging the gap and the model’s true accuracy. After bridging the modality gap, the text data act as better substitutes for image data to evaluate the model’s performance.
E.2 Analysis of the Modality Gap in BLIP and BEiT-3
We expand the VLM Zoo provided by LOVM by adding various variants of BLIP and BEiT-3 to enhance the diversity of the VLM Zoo. We find that the zero-shot image classification performance of the base models of BLIP and BEiT-3 is poor. Therefore, we use the retrieval version of the models provided officially. These models are obtained by fine-tuning the base models on COCO and Flickr30k through contrastive learning, thereby possessing better zero-shot image classification performance. During the pre-training of BLIP and BEiT-3, they may have used multiple loss functions in addition to the contrastive loss (BLIP), or perhaps they did not use contrastive loss at all (BEiT-3). Therefore, the existence of the Modality Gap in the feature spaces of these two models requires further investigation.
We use image samples from the STL-10 dataset and class-related text samples generated by ChatGPT. For each modality, we randomly extract 200 samples and calculate image-to-image, text-to-text, and image-to-text similarities, and plot histograms. We also observe the presence of Modality Gaps in different models through UMAP visualization [36]. Figure 5 and Figure 6 show the experiment result. Through our experiments, we verify that the feature spaces of the BLIP and BEiT-3 retrieval models still exhibit the Modality Gap phenomenon.
Figure 5: The distribution of image-to-image (i2i) cosine similarity, text-to-text (t2t) cosine similarity, and image-to-text (i2t) cosine similarity values for different BEiT-3 and BLIP models.
Figure 6: UMAP visualization of image sample features and text sample features from different BEiT-3 and BLIP models.
E.3 Experiment Result on LOVM’s original VLM Zoo
We have provided experimental results on the VLM Zoo originally provided by LOVM [64]. This VLM Zoo contains 35 VLMs. Table 9 shows the experimental results, and we can see that Swab achieves the best performance across all evaluation metrics.
Table 9: Results on LOVM’s original VLM Zoo. We evaluate our method across 23 datasets and 35 pre-trained VLMs. The results are averaged over all datasets.
Methods H-Score NCE LEEP LogME INB Avg Rank ModelGPT Swab
$R_{5}(\uparrow)$ 0.183 0.235 0.139 0.200 0.443 0.443 $\text{0.446}\scriptstyle\pm\text{0.004}$ $\textbf{0.498}\scriptstyle\pm\text{0.005}$
$\tau(\uparrow)$ 0.000 0.000 0.014 -0.014 0.267 0.261 $\text{0.272}\scriptstyle\pm 0.010$ $\textbf{0.310}\scriptstyle\pm\text{0.012}$
$R_{5}$ + $\tau(\uparrow)$ 0.183 0.235 0.153 0.186 0.710 0.704 $\text{0.718}\scriptstyle\pm\text{0.013}$ $\textbf{0.808}\scriptstyle\pm\text{0.011}$
E.4 Per-Dataset Experiment Result
@@
Table 10: LOVM Benchmark (top-1 accuracy).We compare the per-dataset experiment result (fixing the random seed) between ModelGPT and Swab.
Stanford Cars CIFAR100 CLEVR-DIST. CLEVR-COUNT Country211 Retinopathy DMLab DTD EuroSAT FER2013 Flowers102 GTSRB ImageNet KITTI MNIST PCam Oxford Pets Rendered SST2 RESISC45 STL10 SUN397 SVHN VOC2007 Mean $R_{5}$ ModelGPT 0.80 0.80 0.00 0.40 0.60 0.00 0.00 0.80 0.60 0.40 0.80 0.60 0.80 0.00 0.00 0.00 0.80 0.20 0.80 0.20 0.60 0.60 0.60 0.452 Swab 0.80 0.80 0.00 0.60 0.40 0.20 0.00 0.80 0.60 0.60 0.80 0.60 1.00 0.00 0.40 0.00 0.80 0.20 1.00 0.20 0.60 0.60 0.60 0.504 $\tau$ ModelGPT 0.67 0.67 0.00 0.00 -0.33 0.00 0.00 0.67 1.00 0.00 0.67 1.00 0.67 0.00 0.00 0.00 0.00 0.00 0.67 0.00 -1.00 0.33 1.00 0.261 Swab 0.67 0.91 0.00 0.33 0.00 0.00 0.00 0.55 1.00 0.33 0.33 1.00 0.89 0.00 0.00 0.00 0.33 0.00 0.67 0.00 -0.82 0.82 0.33 0.320